OpenStudy (anonymous):

Find the last two digits of 9^8^7. (By convention, exponent towers are evaluated from the top down, so 9^8^7 = 9^(8^7).)

OpenStudy (agent0smith):

You might want to look for patterns... Try seeing what the last two digits of 9^x are for a few values of x, and see if you can find a pattern. Then use the pattern to help find the last two digits of 9^2097152 (which is 8^7). There might be a pattern in multiples of something, which you can then divide 2097152 by to shortcut.

OpenStudy (anonymous):

oh. oohhh. i see the pattern. 1 then 9 then 1 then 9 forever and ever right? so then since 2097152 is even, then the units digit would be 9. but what about the tens digit?

OpenStudy (anonymous):

wouldn't it be 1 in the units?

OpenStudy (anonymous):

9^2 = 81

OpenStudy (anonymous):

oh yeah. thanks @pgpilot326

OpenStudy (anonymous):

(10 - 1)^ 8^7 use binomial expansion to get at last 2 terms?

OpenStudy (anonymous):

We're interested in determining $$9^{8^7}\pmod 100$$

OpenStudy (anonymous):

oops, that should be \(\pmod{100}\)!

OpenStudy (anonymous):

mod 100 ?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=9^%288^7%29 Last few digits: 081

OpenStudy (anonymous):

wow @eliassaab, that is a really cool "calculator"

OpenStudy (anonymous):

It is much much more than a calculator

OpenStudy (anonymous):

yes, but how...

OpenStudy (agent0smith):

There is a pattern, but it's not all that simple unfortunately... and this is tedious. Last two digits are the right column. ugh. 9^2 = 27 9^3 = 81 9^4 = 65 61 9^5 = 590 49 9^6 = 5314 41 9^7 = 47829 69 9^8 = 430467 21 9^9 = 3874204 89 9^10 = 34867844 01 9^11 = 313810596 09 9^12 = 2824295364 81 9^13 = 25418658283 29 9^14 = 228767924549 61 9^15 = 2058911320946 49 9^16 = 18530201888518 41 9^17 = 166771816996665 69 9^18 = 1500946352969991 21 9^19 = 13508517176729920 89 9^20 = 121576654590569288 01 9^21 = 1094189891315123592 09 9^22 = 9847709021836112328 81 9^23 = 88629381196525010959 29 9^24 = 797664430768725098633 61 9^25 = 7178979876918525887702 49 9^26 = 64610818892266732989322 41 9^27 = 581497370030400596903901 69

OpenStudy (anonymous):

i didn't know what to call it :)

OpenStudy (agent0smith):

2, 8, 6, 4, 4, 6, 2, 8, 0, 0, 8, 2, 6, 4, 4, 6, 2, 8, 0, 0, 8, 2, 6, 4, 4, 6, 2, 8, 0, 0, 8, 2 blech. That doesn't make it very easy.

OpenStudy (anonymous):

you could mod 10 the exponent and then take 9 to the remainder 90^10 = 1 mod 100 => 9^2 = last 2 digits, just as wolfram's calculated

OpenStudy (anonymous):

thanks, everyone, for helping!

OpenStudy (anonymous):

you need the euler totient function to find what power will mod to 1.

OpenStudy (anonymous):

yay! thanks!

OpenStudy (anonymous):

agh! i don't know who to give the medal to so?